Please help me with these questions.

1). Find f(2), f(−7), f(k), and f(k^2 − 1). f(x) = 3/x-5

2). Evaluate the expression.

P(12, 2) · C(12, 2)
Answer #1

Solection: f(x)= 3 3 x-5 f(2)= 3 - 3 = -1 2-5 -3 11 - 3 +(-7)= 3 12 lit 4 -7-5 -12 3. TR) = 3 R-5 3 3 +CR²-2) R²-6 R²-1-5